# Digital currency hardware wallets and the essence of money

### v3

@@ 2.2 Digital currency use cases
One might argue that conducting transactions using digital currency held by custodians is, from the perspective of users and counterparties, not much different than conducting transactions practice of using ordinary bank accounts, with a possible exception of a different regulatory environment for the digital currency custodian. The custodian is still a gatekeeper that could fail to meet its obligations; the user is still bound to the custodian by identification requirements and still does not hold the assets directly. One might also argue that digital currency held by custodians is even less useful than bank deposits, if it is assumed that digital currency tokens cannot be rehypothecated and therefore cannot contribute to money creation.
@@
We might reasonably conclude that non-custodial wallets, which allow users to hold digital assets directly, are a foundational benefit of digital currency. However, even if we accept that non-custodial wallets are essential to digital currency, the idea that offline payments are necessary for digital currency is certainly false. Users have an interest to possess and control digital assets directly, regardless of the context in which they spend them. Holding digital assets directly takes custodians out of outbound transactions, allowing users to avoid the risk that the custodians might extract fees, block certain (or all) transactions from taking place, or build profiles of the behaviour of their customers by monitoring the details of their transactions, such as time, location, and counterparty. Bearer wallets can be used in a variety of contexts, including but not limited to in-person POS transactions. Importantly, the benefits of non-custodial wallets can also be realised in a fully online context such as e-commerce.
Proponents of digital currency are right to acknowledge the secular decline in the use of cash in retail consumer transactions as a primary motivator for digital currency. As fewer consumers use cash for transactions, the variable revenues associated with operating a cash infrastructure are falling below the fixed costs. This pattern is evident in much of the developed world. For example, in the UK, banks are closing ATMs and branches [12], making cash more difficult for consumers to withdraw [13]. Some merchants have decided that the marginal cost of the various expenses related to accepting cash payments, such as tills, employee time, cash accounts with banks, and the means of transporting cash, outweighs the benefit of providing goods and services to cash-only customers. As a result, the value of cash to consumers is reduced, since customers no longer have assurance that their cash will be accepted at the point of sale, creating a feedback loop.
@@
And so, when proponents of digital currency say that digital currency should allow “cash-like” transactions for retail users of money, they are referencing the idea that digital currency can provide individuals with a payment option that provides the key affordances of cash, including direct ownership, privacy, and equal access without discrimination. Users of digital currency would hold direct obligations of the central bank, and, given the right regulations, they might have reason to believe that their digital currency would be accepted by retailers that require electronic transactions.
@@
Digital currency can certainly address problems associated with the systematic decline in the set of available options for consumers to access cash and the increasing tendency for merchants to refuse payments in cash at the point of sale. Users can load non-custodial wallets with the help of a personal device with network access, as well as via facilities provided by banks or ATM networks. Consumers can use non-custodial wallets in-person, to interface with point of sale devices in the same manner that they use debit cards or other custodial payment mechanisms today. Such payments can be considered “offline” both in the sense that they are in-person and in the sense that that the non-custodial wallets themselves do not require Internet connections. We refer to such payment scenarios as single-offline, because only one of the counterparties to a transaction has access to third parties via a network or other means of connectivity. Single-offline payments include in-person EMV card payments via POS systems and are therefore constitute a significant share of the payments that have replaced cash payments in the retail payments sector over the past twenty years.
@@ 2.3 Hold offline, transact online
The assumption that transactions that are not between accounts must be offline is wrong. This assumption would imply that all payments that are not online between payment accounts must be offline between peers that are disconnected from real-time communication with any other party. In particular, the proposal by the European Central Bank (ECB) for the digital euro [14] holds that these two models for payment, online via accounts and dual-offline between peers, are comprehensively exhaustive, but this is simply false. It simply ignores the possibility that money can be held offline and transacted online. Specifically, it is possible that a consumer would hold digital currency tokens in a non-custodial wallet and transfer them to a merchant with online connectivity, in much the same way that a consumer might rely upon the online connectivity of a POS device offered by the merchant.
@@
There is a clear case for using digital currency with non-custodial wallets as a substitute for electronic card payments or network platform transactions at the point of sale. What remains is the question of whether there exists a need to support payments between two parties such that neither of which is connected to a network. There is no evidence that consumers who cannot withdraw cash or merchants who refuse to accept cash are using alternatives to cash that do not make use of a network connection instead, and there is no evidence of a significant demand for one, either. As a means of facilitating a transaction between two parties in situations in which neither party has a networked device, cash is remarkably effective. Although it is possible to imagine narrowly construed use cases for new capabilities, such use cases do not justify system requirements.
@@
It is worth asking whether governments or central banks are harbouring an intention to do away with cash entirely, or whether the presumption of such an intention is shaping the view of what CBDC should be. For example, some authorities have argued that a larger share of domestic transactions should be brought inside the scope of the formal economy [15, 16]. Legitimate or not, that goal is not inconsistent with prevailing mechanisms to roll out digital transactions in general and does not require CBDC. Moreover, few monetary authorities or government regulators have called for the abolition of cash, whereas many, including those of Switzerland, the United Kingdom, and the United States, have pledged to preserve access to cash for the foreseeable future. Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [17]. So the idea that CBDC is mostly a justification for phasing out cash, as has been speculated by some, is not corroborated by facts.
@@ 3 Certified hardware as a de facto custodian
Design approaches based on certified hardware have featured prominently in the recent debate concerning the deployment of large-scale digital currency systems in general and systems to support CBDC in particular. Certified hardware is a cornerstone of the so-called “trusted computing” paradigm, wherein a user carries a device designed to operate in a manner that respects the will of a third party in preference to the will of the user. The justification for such approaches is usually given in terms of the perceived need to facilitate offline payments or to facilitate the recovery of lost assets. In this section, we examine the characteristics of trusted computing within the context of its application to digital currency systems and its potential impact on the power relationships between the users of devices, the manufacturers of devices, and other powerful actors such as corporations and the state.
Baressi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [5], implicitly drawing the hardware manufacturer into the role of trusted third party. In their model, the wallet would be a means of transacting with money held in custodial accounts that would be reconciled ex post.
@@ 3.1 Fair exchange
If we define a fair exchange as a transaction mechanism that ensures that transacting parties either each receive or each do not receive what it expects (and do not receive “any additional information about the other’s item”) [19], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [20]. With the exchange of physical objects, possession of the object itself is sufficient to demonstrate the veracity of the exchange that has taken place. But with digital objects the situation is more complicated. It is possible to have fair exchange without connection to the Internet, provided that there are third parties accessible locally, perhaps via a local network, as might be possible with ad hoc networks.
Proponents of approaches requiring certified hardware have proposed embedding into end-user devices a tamper-resistant “secure element” that can be used to systematically restrict what the device can do, disclose the immutable identity of the device to a third party, or furnish information to third parties about what the device had observed or done. The secure element would be impenetrable to the user, and the user would not be able to modify its behaviour without rendering the device inoperable. In effect, such an element effectively acts as the third party by proxy, by embedding the will of the third party into the operation of a device carried by a user.
@@ 3.2 The perils of trusted hardware
In the case of digital currency, the secure element would presumably be used as an on-board authority that would prevent double-spending by maintaining a record of transactions made by the device and disallowing users from transacting tokens that have already been signed over to others. This design model in particular, and trusted hardware in general, introduces a variety of problems. We consider some of those problems here:
@@
    Trusted computing introduces significant security risks particular to digital currency. The risks of relying upon the security of a hardware manufacturer are too great for a token-based CBDC context. In particular, a system design that relies upon trusted hardware components is fragile by design. If the trusted hardware fails to function as advertised or is compromised by an attacker, then users could potentially transact the same tokens an arbitrary number of times without bound. The system-level costs of such an attack would be huge, and the incentives would be tremendously enticing for attackers. We note that modern devices, such as debit cards, that rely upon tamper-proof hardware for the exchange of value generally interact with accounts. If their security is compromised, then the value available to attackers is limited by rules and operational procedures implemented by custodians. Without custodial accounts, such backup mechanisms are not possible, and so, a successful attack on the hardware root of trust would imply that it would be necessary for all users to replace their devices. Thus, it is reasonable to conclude that it is preferable to rely upon a protocol to achieve the required level of security, as can be done when one of the parties has access to the network, rather than to expect that trusted hardware modules would guarantee that all transactions would take place within the security envelope of a trusted custodian. To assume that all hardware wallets must take the latter approach because some circumstances might require it is tantamount to overspecification at the expense of users who would readily hold assets offline but transact them offline. It is worth noting that when Chaum, Grothoff, and Möser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [34], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online.
@@
    Secure elements are unaccountable to their users. Because trusted hardware is impenetrable to users, users cannot know with certainty what their devices are doing. The inability to analyse the behaviour of a device undermines the trust that users can and should place in the device. In essence, the device serves a second master that is not the user. Users cannot trust what they cannot verify, so they have every reason to be suspicious of unchecked promises on the part of hardware manufacturers or other privileged guarantors [21]. For example, secure elements can furnish information that can be used to identify the user or the user’s previous transactions. By contrast, with open hardware, skilled members of the community can conduct audits and provide reports to the public. Of course, users with the means to do so can personally conduct the audits themselves or hire contractors to conduct the audits for them. However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [22].
@@
    Certified hardware discourages innovation. Users who cannot create their own devices cannot innovate. Instituting a certification requirement for hardware serves to ensure that only those with the time and resources to pursue certification will be able to produce working devices. Regulations that increase the cost of entry are known to decrease entrepreneurship as well as innovation by existing firms [23]. Ordinary users and hobbyists will not be able to develop their own devices, creating a barrier to public understanding the of the core technology, encumbering public debate and consideration of alternative designs [24, 25].
@@
    Certified hardware distorts the market by privileging its manufacturers. The requirement to use trusted hardware effectively locks the hardware manufacturer in to a privileged position in the market, wherein prospective manufacturers require certification by some authority (either directly by the government or by a delegated authority such as a trade organisation), and concomitant lobbying and regulation can restrict free entry, increasing market concentration [26]. Furthermore, even if all manufacturers could have certification at zero marginal cost, the fact that fabrication facilities carry high fixed costs means that the market is likely to be highly concentrated, with manufacturers capturing the surplus [27].
@@
    A certified device is tantamount to a relationship with a custodian. Hardware manufacturers become gatekeepers at best, and globally trusted third parties in the general case. But unlike trusted third parties on the network, users bind themselves to long-term relationships with personal devices that contain trusted hardware; the relationships are not much different between account-holders and custodians, wherein account-holders are subject to satisfying the requirements of custodians. The cost of switching bank accounts is generally only a matter of time and effort on the part of account-holders, whereas the cost of switching devices generally includes per-device manufacturing and retailing cost plus taxes and costs associated with physical acquisition and disposal of the devices. In addition, while it is possible to implement both technical protocols and regulations that rely upon locally-trusted third parties, economies of scale mean that hardware manufacturers are likely to have global reach. It is reasonable to conclude that relationships between users and their devices featuring specific trusted hardware will be sticky indeed.
For these reasons, we are unconvinced that trusted hardware solutions are appropriate as requirements for the design of tools that allow users to possess and control their own assets. At the same time, we would not suggest that trusted hardware should be forbidden. In some fraction of cases in which physical cash is not an option and network connections are not available, trusted hardware might be an alternative kind of third party for transacting parties to consider, although this possibility is not sufficient justification to institute a requirement. In all cases, both sender and recipient must willingly choose to trust the third party to consummate the transaction, and given the final three points in the last above, it is difficult to imagine that a user preference for a particular device manufacturer would be real. Consider, for example, that existing users of secure hardware in the form of electronic payment cards generally have little say in which payment network their issuing banks choose.

### v4

@@ 2.2 Motivations for digital currency
One might argue that conducting transactions using digital currency held by custodians is, from the perspective of users and counterparties, not much different than conducting transactions through ordinary bank accounts, with a possible exception of a different regulatory environment for the digital currency custodian. The custodian is still a gatekeeper that could fail to meet its obligations; the user is still bound to the custodian by identification requirements and still does not hold the assets directly. One might also argue that digital currency is less useful than cash if it is assumed that it can held only by depository institutions, or that individuals cannot deposit and withdraw it from depository institutions, or if it cannot be exchanged by depository institutions for central bank reserves.
@@
We might reasonably conclude that non-custodial wallets, which allow users to hold digital assets directly, are a foundational benefit of digital currency. However, even if we accept that non-custodial wallets are essential to digital currency, the idea digital currency held offline must be spent offline is certainly false. Users have an interest to possess and control digital assets on devices of their choosing, rather than devices associated with attributable accounts, regardless of the context in which they spend them, for reasons of privacy and censorship resistance [Kahn 2005], which we shall address in Section 3. Holding digital assets directly takes custodians out of outbound transactions, allowing users to avoid the risk that the custodians might extract fees, block certain (or all) transactions from taking place, or build profiles of the behaviour of their customers by monitoring the details of their transactions, such as time, location, and counterparty. Bearer wallets can be used in a variety of contexts, including but not limited to in-person POS transactions. Importantly, consumers can realise the privacy and anti-censorship benefits of non-custodial wallets in a fully online context such as e-commerce, for example, if they are using a system that supports asymmetric privacy [Tinn 2024].
Proponents of CBDC are right to acknowledge the secular decline in the use of cash in retail consumer transactions as a primary motivator for digital currency. For example, in the UK, banks are closing ATMs and branches [Access to Cash Review 2019], making cash more difficult for consumers to withdraw and for businesses to manage cash [Tisher 2020]. Although court judgments in some jurisdictions suggest that the legal tender status of banknotes implies mandatory acceptance in all but exceptional circumstances [European Union 2021], jurisprudence in other jurisdictions only recognises legal tender as something that must be accepted in the settlement of a debt [Bank of England 2025], admitting the possibility that vendors can freely choose what forms of payment they will accept in the general case.
@@
And so, when proponents of CBDC say that CBDC should allow “cash-like” transactions for retail users of money, they are foreshadowing the idea that CBDC can provide individuals with a payment option that provides the key affordances of cash, including direct control, privacy, and equal access without discrimination. Users of digital currency would hold direct obligations of the central bank, as they do with banknotes issued by the central bank, and, given the proper regulations, they might have reason to believe that their digital currency would be accepted by retailers that prefer electronic transactions.
@@
CBDC can certainly address problems associated with the systematic decline in the set of available options for consumers to access cash and the increasing tendency for merchants to avoid payments in cash at the point of sale. Users can load non-custodial wallets with the help of a personal device with network access, as well as via facilities provided by banks or ATM networks. Consumers can use non-custodial wallets in-person, to interface with point of sale devices in the same manner that they use debit cards or other custodial payment mechanisms today. Such payments can be considered “offline” both in the sense that they are in-person and in the sense that that the non-custodial wallets themselves do not require Internet connections. We refer to such payment scenarios as single-offline, because only one of the counterparties to a transaction has access to third parties via a network or other means of connectivity. Single-offline payments include in-person EMV card payments via POS systems and therefore constitute a significant share of the payments that have replaced cash payments in the retail payments sector over the past twenty years.
@@ 2.3 Hold offline, transact online
The assumption that transactions that are not between accounts must be offline is wrong. This assumption would imply that all payments that are not online between payment accounts must be offline between peers that are disconnected from real-time communication with any other party. In particular, by stating that “digital euro users should have the choice to use the digital euro either online or offline, or both”, the proposal by the European Central Bank (ECB) for the digital euro [European Commission 2023] implicitly suggests that the two models it describes for payment, online via accounts and dual-offline between peers, are comprehensively exhaustive, but this is simply false. It simply ignores the possibility that money can be held offline and transacted online. Specifically, it is possible that a consumer would hold digital currency tokens in a non-custodial wallet and transfer them to a merchant with online connectivity, in much the same way that a consumer might rely upon the online connectivity of a POS device offered by the merchant.
@@
Given the rise of e-commerce transactions, the tendency of merchants to refuse cash at the point of sale, and the legitimate interest of consumers in privacy and censorship avoidance, there is a clear case for using CBDC with non-custodial wallets as a substitute for electronic card payments or network platform transactions at the point of sale. What remains is the question of whether there exists a need to support payments between two parties such that neither of which is connected to a network. There is no evidence that consumers and merchants who dislike cash are seeking alternative payment technologies that do not make use of a network connection. As a means of facilitating a transaction between two parties in situations in which neither party has a networked device, cash is remarkably effective. Although it is possible to imagine narrowly construed use cases for dual-offline payments, such as transient network outages, those use cases do not justify system-level requirements for a hardware root of trust on all end-user devices.
@@
It is worth asking whether governments or central banks are harbouring an intention to do away with cash entirely, or whether the presumption of such an intention is shaping the view of what CBDC should be. The argument that a larger share of domestic transactions should be brought inside the scope of the formal economy is certainly not new [FATF 1996, Camera 2001]. Recently, the governments of certain countries, including India [Ministry of Finance 2025] and Indonesia [Septiningrum 2024], have taken steps toward eliminating cash entirely, an objective that is consistent with prevailing business-led approaches to phase out cash transactions and encourage digital transactions in general. However, achieving that goal, whether it is legitimate or not, does not require CBDC; some countries have already implemented regulations to prevent the use of cash for large transactions or to control bank deposits comprising large amounts of cash [Financial Conduct Authority 2023]. Moreover, few monetary authorities or government regulators have called for the abolition of cash, whereas many, including those of Switzerland [Reuters 2024], the United Kingdom [Bank of England 2024], and the United States [Federal Reserve 2022], have pledged to preserve access to cash for the foreseeable future. Notably, Sweden has rolled back its drive toward creating a cashless society to support cash infrastructure in a near-cashless economy, at great cost [Library of Congress]. So the idea that CBDC is mostly a justification for phasing out cash, as has been speculated by some, is not corroborated by the facts, not yet at least.
@@ 3 Certified hardware as a de facto custodian
Design approaches based on certified hardware have featured prominently in the recent debate concerning the deployment of large-scale digital currency systems in general and systems to support CBDC in particular. Certified hardware is a cornerstone of the so-called “trusted computing” paradigm, wherein a user carries a device designed to operate in a manner that respects the will of a third party rather than the will of the user. The justification for such approaches is usually given in terms of the perceived need to facilitate offline payments or to facilitate the recovery of lost digital assets (for instance, in case of loss or theft of the hardware device in which those assets were stored). In this section, we examine the characteristics of trusted computing within the context of its application to digital currency systems and its potential impact on the power relationships between the users of devices, the manufacturers of devices, and other powerful actors such as corporations and the state.
Barresi and Zatti explicitly assume the use of trusted hardware to draw transactions into the security envelope of a trusted custodian [Barresi 2023], implicitly drawing the hardware manufacturer into the role of trusted third party. In their model, the wallet would be a means of transacting with money held in custodial accounts that would be reconciled ex post.
@@ 3.1 Fair exchange
If we define a fair exchange as a transaction mechanism that ensures that “at the end of an exchange [between two players], either each player receives the item it expects or neither player receives any additional information about the other’s item” [Asokan 1998], then it is not possible to have a fair exchange between two parties without the involvement of a third party; this is a foundational truth [Pagnia 1999]. With the exchange of (unregistered) physical objects, possession of the object itself can be sufficient to demonstrate the veracity of the exchange that has taken place. But with digital objects the situation is more complicated. It is possible to have fair exchange without connection to the Internet, provided that a mutually trusted third party is accessible locally, perhaps via a local network, as might be possible with ad hoc networks. In particular, it is not necessary for transacting parties to have connectivity to the issuer of the assets being transacted, since the mutually trusted third party might not be the issuer, and it is not necessary for transacting parties to have connectivity to providers of attributable accounts, such as banks, since the mutually trusted third party might not be an account provider or custodian. Transacting parties can provide assurance, for example, in the form of cryptographic proofs or signatures, that any applicable redemption and compliance obligations will be fulfilled.
Proponents of approaches requiring certified hardware have proposed embedding into end-user devices a tamper-resistant “secure element” that can be used to systematically restrict what the device can do. Such restrictions can be used to provide the immutable identity of the device to a third party, or to furnish information contained within the device to third parties [IETF RFC 9334 2023, Bela 2024]. The secure element would be impenetrable to the user, and the user would not be able to modify its behaviour without rendering the device unusable. In effect, such an element effectively acts as the third party by proxy, by embedding the will of the third party into the operation of a device carried by a user.
@@ 3.2 The perils of trusted hardware
In the case of digital currency solutions that rely upon a hardware root of trust, the secure element would presumably be used as an on-board authority that would prevent double-spending by maintaining a record of transactions made by the device and disallowing users from transacting tokens that have already been signed over to others. This design model in particular, and trusted hardware in general, introduces a variety of problems that can be avoided if support for dual-offline payments is not mandatory. We consider some of those problems here:
@@
    Trusted computing introduces significant security risks particular to digital currency. The risks of relying upon the security of a hardware manufacturer are too great for a token-based CBDC context. In particular, a system design that relies upon trusted hardware components is fragile by design. If the trusted hardware fails to function as advertised or is compromised by an attacker, then users could potentially transact the same tokens an arbitrary number of times without bound. Whereas with account-based systems, such as credit and debit cards, the cost associated with the compromise of a hardware device is limited to the size of the account with which it is associated, for a privacy-respecting system in which money is not linked to identity of the payer, the system-level costs of such an attack would be huge, and the incentives would be tremendously enticing for attackers. We note that modern devices, such as debit cards, that rely upon tamper-proof hardware for the exchange of value generally interact with accounts [ISO 8583 2023]. If their security is compromised, then the value available to attackers is limited by rules and operational procedures implemented by custodians. Without custodial accounts, such backup mechanisms are not possible, and so, the consequences of a successful attack on the hardware root of trust, such as one that impacts the cryptographic algorithm implemented in hardware or the mechanism that protects it from tampering, might not be addressable by a software upgrade, in turn requiring all users to replace their devices. Therefore, if both privacy and security are objectives of a digital currency system, it is preferable to rely upon a protocol, rather than the impermeability of end-user devices, to enforce the rules necessary to achieve the required level of security. Such a protocol can be used whenever one of the parties has access to the network, and under such circumstances, it would be better to allow the use of such a protocol rather than assume that trusted hardware modules would not be breached and would guarantee that all devices operate within the security envelope of a trusted custodian. To assume that all hardware wallets must take the latter approach because some circumstances might require it is tantamount to overspecification at the expense of users who would readily hold assets offline but transact them online. It is worth noting that when Chaum, Grothoff, and Moser argue that “any economically producible device that stores tokens with monetary value in an individual’s possession, and enables offline transactions… will be the target of successful forgery attacks…” [Chaum 2021], they are referring to wallets that perform dual-offline transactions, in contrast to wallets that hold tokens offline and transact them online. While overspecification of device requirements might certainly appeal to device manufacturers seeking compulsory mandates for their proprietary technology, it is important to consider that protocols that do not rely upon a hardware root of trust can function perfectly well outside the dual-offline context. Such protocols do not depend upon the imperviousness of end-user devices to attacks by those who possess them, and such attacks tend to fall within the capabilities of adversarial state actors and global crime syndicates.
@@
    Secure elements are unaccountable to their users. Because trusted hardware is impenetrable to users, users cannot know with certainty what their devices are doing. The inability to analyse the operations of a device undermines the trust that users can and should place in the device. In essence, the device serves a second master that is not the user. Users cannot trust what they cannot verify, so they have reason to be suspicious of unchecked promises on the part of hardware manufacturers or other privileged guarantors [Stallman 2002]. For example, tamper-proof components might store and furnish information that can be used to identify the user or the user’s previous transactions. By contrast, with free hardware design111“design that permits users to use the design (i.e., fabricate hardware from it) and to copy and redistribute it, with or without changes” [Stallman 2015]., skilled members of the community, unbound by nondisclosure agreements, can conduct audits on the design of the hardware and provide reports to the public. Of course, users with the means to do so can personally conduct the audits themselves or hire contractors to conduct the audits for them. However, it is not necessary for every user to actually run an audit; evidence suggests that similar benefits are achieved so long as all users can perform audits if they choose to do so [Heydt-Benjamin 2006].
@@
    Certified hardware discourages beneficial innovation. Users who cannot create their own devices cannot innovate. Instituting a certification requirement for hardware serves to ensure that only those with the time and resources to pursue certification will be able to produce working devices. Regulations that increase the cost of entry are known to decrease entrepreneurship as well as innovation by existing firms [Kozeniauskas 2022]. Ordinary users and hobbyists will not be able to develop their own devices, creating a barrier to public understanding the of the core technology, encumbering public debate and consideration of alternative designs [Felten 2002, Samuelson 2016]. For this reason, hardware certification requirements may implicitly promote security through obscurity, contrary to accepted practices in security engineering [Anderson 2001].
@@
    Certified hardware undermines competition by privileging its manufacturers. The requirement to use trusted hardware effectively locks the hardware manufacturer in to a privileged position in the market, wherein prospective manufacturers require certification by some authority (either directly by the government or by a delegated authority such as a trade organisation), and concomitant lobbying and regulation can restrict free entry, increasing market concentration [Gutierrez and Philippon 2019]. In essence, enforcing hardware certification transforms a market for consumer devices into something more closely resembling public infrastructure procurement. Furthermore, even if certification were not required or if all manufacturers could have certification at zero marginal cost, the fact that fabrication facilities carry high fixed costs means that the market is likely to be highly concentrated, with a few dominant manufacturers capturing the surplus [Gao and Kehrig 2017]. Without a market to serve their needs directly, end-users will have no choice but to accept the solution that the government chooses to trust.
@@
    A certified device is tantamount to a relationship with a custodian. Hardware manufacturers become gatekeepers at best, and globally trusted third parties in the general case. But unlike trusted third parties on the network, users bind themselves to long-term relationships with personal devices that contain trusted hardware; the relationships are not much different between account-holders and custodians, wherein account-holders are subject to satisfying the requirements of custodians. Although in principle, this can be regulated, implementation inexorably depends upon how custodians implement regulatory compliance and is subject to the limits of the effectiveness of regulation. The cost of switching bank accounts is generally only a matter of time and effort on the part of account-holders and nevertheless remains a notoriously difficult problem for regulators, whereas the cost of switching devices, such as smartphones, generally includes per-device manufacturing and retailing cost plus taxes, as well as costs associated with physical acquisition and disposal of the devices. In addition, while it is possible to implement both technical protocols and regulations that rely upon locally-trusted third parties, economies of scale mean that hardware manufacturers are likely to have global reach. It is reasonable to conclude that relationships between users and their devices featuring specific trusted hardware will be sticky indeed.
For these reasons, we are unconvinced that trusted hardware solutions are appropriate as requirements for the design of tools that allow users to possess and control their own digital assets. At the same time, we would not suggest that trusted hardware should be forbidden. In some fraction of cases in which physical cash is not an option and network connections are not available, trusted hardware might be an alternative kind of third party instrument for transacting parties to consider, although this possibility is not sufficient justification to institute trusted hardware as a universal requirement. In all cases, both sender and recipient must willingly choose to trust the third party to consummate the transaction, and given the final three points in the list above, it is difficult to imagine that users would be able to exercise free choice of whom to trust rather than be forced to accept a particular hardware manufacturer. Consider, for example, that existing users of secure hardware in the form of electronic payment cards often have little say in which four-point payment network their issuing banks choose, and even when they do, the set of existing card payment networks remains well-entrenched. Similarly, secure hardware solutions have implications for interoperability: If the payer and payee must use the same third-party hardware manufacturer, then network effects might tend to entrench a single platform or small set of platforms.